# AllSpark: a multimodal spatiotemporal general model

### 31 Dec 2023 v1

@@ 2.2 Multi-modal Model
While the single-modal methods in Section 2.1 have demonstrated excellent performance within their respective modalities, they often face challenges in generalizing across multiple modalities due to their construction based on specific prior assumptions. Recognizing the intrinsic requirement for intelligent models to utilize multimodal data for geographic object cognition, numerous researchers have endeavored to balance the cohesion and autonomy among modalities to build multimodal models.
@@
For joint interpretation of two modalities, H. Li and X.-J. Wu recognized the enhancement possible through the fusion of RGB and Infrared images, which convey information about color, texture, and temperature. Given the widespread demand and application prospects in military, medical, and other fields, they proposed a dense fusion method for RGB and Infrared images: DenseFuseli2018densefuse . This method, employing a dense module that connects the outputs of all convolutional layers, successfully achieves pixel-level fusion of RGB and Infrared modalities. A. Sadeghian et al. extended the use of historical trajectories for predicting multiple plausible trajectories in Social GANgupta2018social . They introduced RGB images to enhance scene data in Sophiesadeghian2019sophie , achieving better results in trajectory prediction tasks. Recognizing the high complementarity between RGB and SAR images, L. H. Hughes et al. proposed a three-step deep neural network framework that utilizes a universal prediction of matching regions, generates heatmaps, and eliminates outliers to match RGB and SAR imageshughes2020deep . X. Li et al. introduced DTCDNli2021deep , a model that employs a GAN network to migrate RGB and SAR images to the same feature space, facilitating target detection. J. Yang et al. proposed a dual-stream convolutional network that uses high-resolution multispectral images to enhance the spatial resolution of hyperspectral imagesyang2018hyperspectralandmultispectral .
@@
Notably, CLIPradford2021learning proposed by A. Radford et al. associates the RGB modality with the text modality using contrastive learning. With pretraining guided by weak supervisory signals from text, CLIP demonstrates outstanding capabilities in both visual single-modal tasks and visual-text multimodal tasks, inspiring a series of subsequent workslu2019vilbert ; kim2021vilt ; alayrac2022flamingo ; li2022blip . Y. Zhang et al. introduced Meta-transformerzhang2023meta , leveraging the contrastive learning paradigm from CLIP to pretrain a universal backbone network under the visual-language modality. It exhibits multimodal generalization abilities across various modalities such as point clouds, infrared, and hyperspectral. J. Han et al. directly employed a multimodal large language model as a universal backbone network, proposing the One-LLM, successfully unifying eight modalities including Image, Audio, Video, and Pointhan2023onellm . The success of these approaches imply the unique role of language modalities in multimodal model.
@@
Therefore, building upon the aforementioned efforts, we systematically propose the fundamental principle, Language as Reference Framework (LaRF). Guided by this principle, we balance the cohesion and autonomy among diverse modalities and introduce a general intelligent model named AllSpark, which unifies 13 spatiotemporal modalities and possesses the potential to extend to an arbitrary number of modalities.
3 Method
3.1 Theory
@@
We observe that in the process of comprehending the world, humans integrate information from multiple modalities such as hearing, touch, smell, and vision. The concepts formed through the parsing of these modalities are ultimately converged in language. It is through language that humans engage in associating, reasoning, and expressive behaviors. In other words, language precisely encodes human perception and understanding of the world, providing clear definitions and meanings to abstract concepts from each modality.
@@
Inspired by our observation, to balance the cohesion and autonomy among multiple modalities, we introduce the fundamental principle, Language as Reference Framework (LaRF). The significance is:
@@
    Alignment Capability: Language can accurately encode both cohesion and autonomy information across multiple modalities. Aligning each modality with the language modality enables the attainment of a unified representation in the same feature space, addressing the challenge of high heterogeneity among modalities.
@@
    Reasoning Capability: Language, being a tool for human thought and expression, inherently possesses the ability to perform complex reasoning. Each modality, when represented in a unified space with LaRF, inherits the reasoning capability of language, unlocking the potential for multimodal joint reasoning.
@@
    Interpretability: Deep learning have long been characterized as opaque ”black boxes.” However, a multimodal intelligent system constructed based on the LaRF can directly leverage language as a tool. This facilitates the output of interpretable reasoning chains that humans can understand, thereby achieving true explainable artificial intelligence.
@@
    Scalability: The multimodal system guided by the LaRF is agnostic to specific modalities. New modalities only need to establish a mapping to the language model to participate in joint reasoning with other modalities. Therefore, theoretically, a multimodal model based on LaRF can be extended to an arbitrary number of modalities.
In summary, we contend that intelligent models, when engaged in multimodal perception processes akin to those of humans, should adopt language as the reference framework for abstract concepts across modalities. The analysis and reasoning should occur in the feature space of language. As illustrated in the conceptual model in Figure 2, when constructing specific multimodal models, data from each modality should be encoded into modality-independent tokens with contextual information. These modality-specific tokens should align with the tokens of language. A universal intelligent model, primarily based on the Transformer architecture, can then be employed for parsing and reasoning, ultimately generating a comprehensive multimodal representation.
Figure 2: The conceptual model with LaRF
3.2 Overview
Table 1: List of symbols
@@ 3.2 Overview
$f_{i}$ modal encoder for $q$ query vectors of the bridge
$\Phi$ modal bridge $s_{i}$ tokens of $m_{i}$
T text tokenizer $p_{i}$ text prompt for $m_{i}$
@@
Multi-modal LLM
$\theta$
model parameters
@@
$y$
label
$L$ loss function $t_{i}$ tokens of $m_{i}$
@@
$W$
weights of the linear layer
@@
$w_{i}$
word of the code or text modal
@@
Guided by the principles mentioned above, the AllSpark we proposed consists of five modules: Modal Encoder, Modal Bridge, Text Tokenizer, Multimodal Large Language Model, and Task Head. The overall architecture is depicted in Figure 3.
Figure 3: AllSpark Architecture
To maintain the autonomy among modalities, we designed modality-specific encoders to encode highly heterogeneous data into modality-independent tokens (Section 3.3 for details). However, the dimens ion of tokens outputted by modal encoders are still inconsistent. To achieve the goal of parsing in the unified representational space of language, we introduced the modal bridge from the Lynxzeng2023matters . The modal bridge aims to project tokens from each modality into the dimension of the Multimodal Large Language Model (Section 3.4 for details). The formalization of this process is as follows:
@@
Here, $\{m_{RGB},m_{MSI},m_{HSI},\ldots,m_{i},\ldots,m_{Video}\}$ represents inputs from various modalities, where $m_{RGB}$ represents the RGB modality, $m_{MSI}$ represents the MSI modality, and so on. $f_{i}$ is defined as the modal encoder for the $m_{i}$ modality, $\Phi$ represents the modal bridge, $q\in R^{N*D}$represents $N$ learnable vectors of dimension $D$ in the modality bridge, where $D$ is set to 4096, representing the dimensionality of the Multimodal Large Language Model. Each modality’s input data $m_{i}$ is mapped to a token sequence $s_{i}\in R^{N*D}$ of the same dimension as the language model. Table 1 summarizes the main mathematical symbols and their meanings in the model.
To achieve cohesion among modalities, we employ a unified Multimodal Large Language Model to parse data from various modalities. The text tokenizer and Multimodal Large Language Model in AllSpark are based on the visual-language model Lynxzeng2023matters . To extend Lynx to 13 spatiotemporal modalities, we design specific text prompts for each modality to guide the model in correctly parsing information from each modality. Additionally, Lynx incorporates several lightweight multimodal adapter layers internally to accommodate multimodal inputs. We continue this design and do not freeze the parameters of the adapter layers during training to enhance the model’s adaptability to other spatiotemporal modalities. Finally, we acknowledge that there exists a gap between the parsing results of the model and the downstream task. Therefore, we design specific task heads for each task to enhance the model’s generalization capability.
@@
All tasks in our experiments are supervised tasks, with $y$ denoting the labels, $L$ representing the loss function, and $\theta$ representing the learnable parameters in our model. The optimization objective of the model can be formalized as:
$\theta_{i}=\mathop{\arg\min}_{\theta}{L\left(y,M\left(m_{i},p_{i}\right),\theta_{i}\right)}$ (3)
3.3 Independent Encoder for Each Modality
@@
The modal encoder aims to encode the raw data of each modality into a token sequence, formalized as: $t_{i}=f_{i}\left(m_{i}\right)$, where $t_{i}\in R^{n*d}$. We have designed different modal encoders for each modality to maintain autonomy among modalities. The following provides individual introductions for each modality:
3.3.1 1D modal
@@
Code & Text: Code is essentially a specialized form of language, but due to its distinct properties—such as having a strict syntax, being unambiguous, and being capable of interacting with machines—we separate it as a distinct modality. Exploring intelligent methods for the Code modality is crucial for reliable AI reasoning and achieving interaction between intelligent models and the real world. Since the Lynx model is a language model, we do not design an additional modal encoder for the Code and Text modalities. Instead, we directly utilize Lynx’s text tokenizer, i.e., $f\left(m_{Text/Code}\right)=T\left(m_{Text/Code}\right)$, where $m_{Text/Code}=\{w_{1},w_{2},w_{3},\ldots\}$ represents the sequence of words in text or code.
3.3.2 2D modal
@@
RGB: RGB imagery represents the visible light spectrum, reflecting the electromagnetic characteristics of objects that emit or reflect visible light waves. It is the most common modality in visual research. RGB imagery is a standard three-band image, i.e., $m_{RGB}\in R^{H*W*3}$. For the modal encoder of this modality, we adopted the visual encoder from the Lynx model: EVA. EVA is a large visual backbone model composed of 40 stacked Transformer Blocks with a width of 1408. During the experiments, AllSpark loaded the official weights of the EVA model and froze them for training.
@@
MSI: Multispectral imagery is a modality extensively studied in the remote sensing field. Typically, it includes more than the three visible light bands present in the RGB modality. MSI incorporates multiple non-visible light bands such as near-infrared, short-wave infrared, coastal atmospheric aerosol, and cirrus bands. Therefore, the number of channels in the imagery is usually greater than the three bands in RGB imagery, i.e., $m_{MSI}\in R^{H*W*C}$, where $C>3$. We extended the PatchEmbed of the standard Vision Transformerdosovitskiy2020image , modifying its channel count to match the number of bands in the input multispectral imagery. This adaptation allows it to serve as the feature encoder for the multispectral modality.
@@
HSI: Hyperspectral imagery increases the number of bands compared to multispectral imagery, often reaching hundreds of bands, with each band containing rich information about the terrain. Unlike RGB and multispectral imagery, where an image serves as a single sample, in hyperspectral imagery, all bands for each pixel are treated as a single sample, i.e., $m_{HSI}\in R^{1*1*C}$. In the modal encoder for hyperspectral imagery, we first use a linear projection layer to expand the feature dimensions for each pixel. This process is formalized as $W*m_{HSI}^{T}$, where $W\in R^{1*d}$ is the weight matrix of the linear projection layer. Subsequently, we use a 12-layer standard Transformer encoder to extract its features. The entire process can be represented as $f\left(m_{HSI}\right)={Enc}_{12}\left(W*m_{HSI}^{T}\right)$.
@@
Table: The Table modality consists of several rows and columns, where each row represents a record, and each column represents a different attribute, i.e., $m_{Table}\in R^{{row}*{col}}$. The modal encoder for the Table modality inherits the design from TabFormerpadhi2021tabular : first, based on the different degrees of discreteness for each column attribute, we use independent Embedding Layers to encode discrete and continuous values separately. Subsequently, we employ a single-layer Transformer Encoder to further extract features. The entire process can be formalized as $f\left(m_{Table}\right)={Enc}_{1}\left(Emb(m_{Table})\right)$.
@@
Trajectory: The trajectory modality reflects the changing information of an object over time and space, composed of a series of two-dimensional coordinate points, i.e., $m_{Trajectory}\in R^{l*2}$, where $l$ represents the sequence length of trajectory points. The encoder for the trajectory modality inherits the design from TUTRshi2023trajectory : first, a linear layer is used to expand the dimensions of the two-dimensional trajectory features. This step can be formalized as ${W*m}_{Trajectory}$, where $W\in R^{d*\left(l*2\right)}$ is the weight matrix of the linear projection layer. Then, we use a 2-layer Transformer Encoder to extract its features. The entire process can be formalized as $f\left(m_{Trajectory}\right)={Enc}_{2}\left({W*m}_{Trajectory}\right)$.
@@
SAR: The Synthetic Aperture Radar (SAR) modality is a type of active remote sensing that reflects the electromagnetic characteristics of objects with respect to microwave backscatter. Due to different polarization modes, the final product of SAR imagery is typically a two-band or single-band image, i.e., $m_{SAR}\in R^{H*W*2}$. Therefore, we designed a simple three-layer convolutional network as its modal encoder.
@@
Infrared: The dataset we selected for infrared is SYSU-MM01, which is, in fact, a multi-modal re-identification dataset comprising both visible and near-infrared modalities. The model needs to simultaneously process infrared and visible light images, i.e., $m_{Infrared}=\{{I_{r}\in R}^{H*W*3},{I_{i}\in R}^{H*W*3}\}$. We inherited and slightly modified the design of DDAGye2020dynamic , discarding the shared ResNet-50 layers’ parameters in DDAG and employing two independently parameterized ResNet-50 networks. These networks extract features from infrared and visible light images, and this process can be formalized as $f\left(m_{Infrared}\right)={Enc}_{resnet}\left(I_{r}\right)\oplus\ {Enc}_{resnet}\left(I_{i}\right)$.
@@
Graph: A graph is constructed from a series of nodes and edges, where the attributes of the nodes and the adjacency attributes of the nodes reflect the majority of the graph’s features, i.e., $m_{Graph}\in R^{K*d}$, where K is the number of nodes, and d is the node feature dimension. In AllSpark, the graph’s modality encoder is based on STAEformerliu2023spatio , with the main design idea being to first use a linear layer to extend its feature dimension and then use several Embedding Layers to separately encode features such as the node’s own characteristics, spatial characteristics, and temporal characteristics in the graph. The entire process can be formalized as $f\left(m_{Graph}\right)={Emb}_{node}\left({W*m}_{Graph}^{T}\right)\oplus{Emb}_{spatial}\left({W*m}_{Graph}^{T}\right)\oplus{Emb}_{time}\left({W*m}_{Graph}^{T}\right)$, where $W\in R^{{hidden}*d}$ is the weight of the linear layer.
@@
Oblique Photography: The oblique photography modality consists of images from multiple views, i.e., $m_{video}=\{V_{1}\in R^{C*H*W},\ V_{3}\in R^{C*H*W},V_{3}\in R^{C*H*W},\ldots\ ,V_{n}\in R^{C*H*W}\}$, where $V_{i}$ represents the image of each view, and n is the number of views. In the encoder for oblique photography modality, we use the standard ViT model to extract features from multiple view images and concatenate them. The process can be formalized as $f\left(m_{video}\right)={Enc}_{12}\left(V_{1}\right)\oplus{Enc}_{12}\left(V_{2}\right)\oplus{Enc}_{12}\left(V_{3}\right)\oplus\ldots\oplus{Enc}_{12}\left(V_{n}\right)$.
3.3.3 3D modal
@@
Video: The video modality can be considered as the stacking of images in the temporal dimension, i.e., $m_{video}\in R^{T*C*H*W}$. The significant characteristic of the video modality is its high redundancy in both temporal and spatial dimensions. To address this issue, the encoder for the video modality in AllSpark is based on TubeViTpiergiovanni2023rethinking , which performs sparse sampling on the highly redundant data in both time and space to reduce redundancy. The features obtained from sparse sampling are then fed into a 6-layer Transformer Encoder for further feature extraction. The process can be formalized as $f\left(m_{video}\right)={Enc}_{6}\left(Sparse(m_{video})\right)$.
@@
PointCloud: PointCloud are typically composed of three-dimensional coordinates and their feature values, $m_{pointCloud}\in R^{K*\left(d+3\right)}$, where K represents the number of three-dimensional points, and d represents the dimensionality of the point cloud features, reflecting information such as the spatial position, shape, color, and texture of objects. The encoder for the point cloud modality inherits the design from PointBERTyu2022point : firstly, point cloud data is grouped and encoded to unify the number of points simultaneously inputted. This step can be represented as $PointGroup=Grouper\left(m_{pointCloud}\right)\in R^{G*N*3}$, where $G$ represents the number of groups, and $N$ represents the number of points in each group. Next, the grouped results are inputted into a one-dimensional convolutional layer to extract feature vectors for each group:$f_{Group}=Conv1d\left(PointGroup\right)\in R^{G*d}$. Finally, the feature vectors for each group are inputted into a standard 12-layer Transformer Encoder to extract their global features. The entire process can be formalized as $f\left(m_{pointCloud}\right)=Enc_{12}\left(Conv1d(Grouper(m_{pointCloud}))\right)$.
3.4 Modal2Language Bridge

### 25 Feb 2024 v2

@@ II-B Multimodal Model
While the single-modal methods in Section 2.1 have demonstrated excellent performance within their respective modalities, they often face challenges in generalizing across multiple modalities due to their construction based on specific prior assumptions. Recognizing the intrinsic requirement for intelligent models to utilize multimodal data for geographic object cognition, numerous researchers have endeavoured to balance the cohesion and autonomy among modalities to construct multimodal models.
@@
For joint interpretation of two modalities, H. Li and X.-J. Wu recognized the enhancement possible through fusing RGB and infrared images, which convey information about colour, texture, and temperature. Given the widespread demand and application prospects in military, medical, and other fields, they proposed a dense fusion method for RGB and infrared images: DenseFuse[29]. This method, employing a dense module that connects the outputs of all convolutional layers, successfully achieves pixel-level fusion of the RGB and infrared modalities. A. Sadeghian et al. extended the social GANs[25] and introduced RGB images to enhance scene data in Sophie[30], achieving better results in trajectory prediction tasks. Recognizing the high complementarity between RGB and SAR images, L. H. Hughes et al. proposed a three-step deep neural network framework that utilizes a universal prediction of matching regions, generates heatmaps, and eliminates outliers to match RGB and SAR images[31]. X. Li et al. introduced the DTCDN[32], a model that employs a GAN network to migrate RGB and SAR images to the same feature space, facilitating target detection. J. Yang et al. proposed a dual-stream convolutional network that uses high-resolution multispectral images to enhance the spatial resolution of hyperspectral images[33].
@@
Notably, CLIP[5], proposed by A. Radford et al., associates the RGB modality with the text modality using contrastive learning. With pretraining guided by weak supervisory signals from text, CLIP has demonstrated outstanding capabilities in both visual single-modal tasks and visual-language multimodal tasks, inspiring a series of subsequent works[6, 7, 8, 9]. Y. Zhang et al. introduced a meta-transformer[34], leveraging the contrastive learning paradigm from CLIP to pretrain a universal backbone network under the visual-language modality. It exhibits multimodal generalization abilities across various modalities, such as point clouds, infrared, and hyperspectral data. J. Han et al. directly employed a multimodal large language model as a universal backbone network, proposing the One-LLM, which successfully unifies eight modalities, namely, images, audio, videos, and points[35]. The success of these approaches implies the unique role of language modalities in multimodal models.
@@
Therefore, building upon the aforementioned efforts, we systematically propose the fundamental principle of the LaRF. Guided by this principle, we balance cohesion and autonomy among diverse modalities and introduce a general intelligent model named AllSpark, which unifies thirteen spatio-temporal modalities and possesses the potential to extend to an arbitrary number of modalities.
III Method
III-A Theory
@@
We observe that in the process of comprehending the world, humans integrate information from multiple modalities, such as hearing, touch, smell, and vision. The concepts formed through the parsing of these modalities ultimately converge in language. Humans engage in associating, reasoning, and expressive behaviours through language. In other words, language precisely encodes human perception and understanding of the world, providing clear definitions and meanings to abstract concepts from each modality.
@@
Inspired by our observation, to balance cohesion and autonomy among multiple modalities, we introduce the fundamental principle of Language as Reference Framework (LaRF).
Figure 2: Guided by the LaRF principle, multimodal data are transformed into a token-context structure similar to language under their respective prior assumptions while maintaining the autonomy of the multimodalities and achieving cohesion among modalities, thus enabling interpretation of multimodal data in the unified language representation space.
In terms of the cohesion of multimodalities, the high heterogeneity between multiple spatio-temporal modalities is a major challenge, while the LaRF principle defines the alignment anchor between multimodalities as language explicitly. As shown in Figure 2, we observe that language is encoded by tokens and their contexts, and this structure can be extended to most spatio-temporal modalities. Therefore, we can align highly heterogeneous spatio-temporal modalities to language modalities in structure and semantics, enabling multimodal interpretation in the unified representation space of language. In addition, the important guiding role of natural language prompts is also one of the reasons why the LaRF principle can achieve cohesion among modalities.
In contrast, we can independently encode multiple spatio-temporal modalities into token sequences under their respective prior assumptions, so the LaRF principle does not lead to the loss of modal autonomy. More importantly, LaRF is not dependent on specific modalities; therefore, theoretically, as long as token representations of modalities can be obtained, the multimodal model guided by LaRF can be extended to arbitrary modalities.
In summary, the significance of LaRF is as follows:
@@
    Alignment Capability: Language can accurately encode both cohesion and autonomy information across multiple modalities. Aligning each modality with the language modality enables a unified representation in the same feature space, addressing the challenge of high heterogeneity among modalities.
@@
    Reasoning Capability: Language, as a tool for human thought and expression, inherently possesses the ability to perform complex reasoning. Each modality, when represented in a unified space with LaRF, inherits the reasoning capability of language, unlocking the potential for multimodal joint reasoning.
@@
    Interpretability: Deep learning methods have long been characterized as opaque ”black boxes.” However, a multimodal intelligent system constructed based on the LaRF can directly leverage language as a tool. This facilitates the output of interpretable reasoning chains that humans can understand, thereby achieving true explainable artificial intelligence.
@@
    Scalability: The multimodal system guided by the LaRF is agnostic to specific modalities. New modalities need to establish a mapping to the language model only to participate in joint reasoning with other modalities. Therefore, theoretically, a multimodal model based on LaRF can be extended to an arbitrary number of modalities.
Figure 3: AllSpark Architecture. Multimodal data are extracted by their respective modal encoders into token sequences. After undergoing dimension alignment with modal-specific text prompt tokens through a modal bridge, the text prompt tokens and modal tokens enter a large language multimodal model for interpretation. Finally, the interpretation result is matched with downstream tasks by task-specific task heads.
III-B Overview
TABLE I: List of symbols
@@ III-B Overview
$f_{i}$
modal encoder for $m_{i}$
$q$ query vectors of the bridge
$\Phi$
modal bridge
$s_{i}$ tokens of $m_{i}$
$T$
text tokenizer
$p_{i}$ text prompt for $m_{i}$
@@
Multimodal LLM
$\theta$ model parameters
@@
$y$ label
$L$
loss function
$t_{i}$ tokens of $m_{i}$
@@
$W$ weights of the linear layer
@@
$w_{i}$ word of the code or text
@@
$\sigma$
softmax
Guided by the principles mentioned above, the AllSpark we proposed consists of five modules: the modal encoder, modal bridge, text tokenizer, multimodal large language model, and task head. The overall architecture is depicted in Figure 3.
To maintain autonomy among modalities, we designed modality-specific encoders to encode highly heterogeneous data into modality-independent tokens (for details, see Section III-C). However, the dimensions of tokens outputted by modal encoders are still inconsistent. To parse in the unified representational space of language, we introduced the modal bridge from the Lynx[10]. The modal bridge aims to project tokens from each modality into the dimension of the multimodal large language model (Section III-D for details). The formalization of this process is as follows:
@@
Here, $\{m_{RGB},m_{MSI},m_{HSI},\ldots,m_{i},\ldots,m_{Video}\}$ represents inputs from various modalities, where $m_{RGB}$ represents the RGB modality, $m_{MSI}$ represents the MSI modality, and so on. $f_{i}$ is defined as the modal encoder for the $m_{i}$ modality, $\Phi$ represents the modal bridge, and $q\in R^{N*D}$represents $N$ learnable vectors of dimension $D$ in the modality bridge, where $D$ is set to 4096, representing the dimensionality of the multimodal large language model. The input data $m_{i}$ of each modality are mapped to a token sequence $s_{i}\in R^{N*D}$ of the same dimension as the language model. Table I summarizes the main mathematical symbols and their meanings in the model.
To achieve cohesion among modalities, we employ a unified multimodal large language model to parse data from various modalities. The text tokenizer and multimodal large language model in AllSpark are based on the visual-language model Lynx[10]. To extend Lynx to thirteen spatio-temporal modalities, we designed specific text prompts for each modality to guide the model in correctly parsing information from each modality. Additionally, Lynx incorporates several lightweight multimodal adapter layers internally to accommodate multimodal inputs. We continue this design and do not freeze the parameters of the adapter layers during training to enhance the model’s adaptability to other spatio-temporal modalities. Finally, we acknowledge that a gap exists between the parsing results of the model and those of the downstream task. Therefore, we design specific task heads for each task to enhance the model’s generalization capability.
@@
All the tasks in our experiments are supervised tasks, with $y$ denoting the labels, $L$ representing the loss function, and $\theta$ representing the learnable parameters in our model. The optimization objective of the model can be formalized as follows:
$\theta_{i}=\mathop{\arg\min}_{\theta}{L\left(y,M\left(m_{i},p_{i}\right);\theta_{i}\right)}$ (3)
III-C Independent Encoder for Each Modality
@@
The modal encoder aims to encode the raw data of each modality into a token sequence, formalized as $t_{i}=f_{i}\left(m_{i}\right)$, where $t_{i}\in R^{n*d}$. We designed different modal encoders for each modality to maintain autonomy among modalities. The following provides individual introductions for each modality:
III-C1 1D modal
@@
Code & Text: Code is essentially a specialized form of language, but due to its distinct properties such as having a strict syntax, being unambiguous, and being capable of interacting with machines. we separate it as a distinct modality. Exploring intelligent methods for the code modality is crucial for reliable AI reasoning and achieving interaction between intelligent models and the real world. Since the Lynx model is a language model, we do not design an additional modal encoder for the code and text modalities. Instead, we directly utilize Lynx’s text tokenizer, i.e., $f\left(m_{Text/Code}\right)=T\left(m_{Text/Code}\right)$, where $m_{Text/Code}=\{w_{1},w_{2},w_{3},\ldots\}$ represents the sequence of words in the text or code.
III-C2 2D modal
@@
RGB: RGB imagery represents the visible light spectrum and reflects the electromagnetic characteristics of objects that emit or reflect visible light waves. It is the most common modality used in visual research. RGB imagery is a standard three-band image, i.e., $m_{RGB}\in R^{H*W*3}$. For the modal encoder of this modality, we adopted the visual encoder from the Lynx model: the EVA. The EVA is a large visual backbone model composed of 40 stacked transformer blocks with a width of 1408. During the experiments, AllSpark loaded the official weights of the EVA model and froze them during training.
@@
MSI: Multispectral imagery is a modality extensively studied in the remote sensing field. Typically, it includes more than the three visible light bands present in the RGB modality. MSIs incorporate multiple nonvisible light bands, such as near-infrared, shortwave infrared, coastal atmospheric aerosol, and cirrus bands. Therefore, the number of channels in the imagery is usually greater than the three bands in RGB imagery, i.e., $m_{MSI}\in R^{H*W*C}$, where $C>3$. We extended the PatchEmbed of the standard ViT[18], modifying its channel count to match the number of bands in the input multispectral imagery. This adaptation allows it to serve as the feature encoder for the multispectral modality.
@@
HSI: Hyperspectral imagery increases the number of bands compared to multispectral imagery, often reaching hundreds of bands, with each band containing rich information. Unlike RGB and multispectral imagery, where an image serves as a single sample, in hyperspectral imagery, all bands for each pixel are treated as a single sample, i.e., $m_{HSI}\in R^{1*1*C}$. In the modal encoder for hyperspectral imagery, we first use a linear projection layer to expand the feature dimensions for each pixel. This process is formalized as $W*m_{HSI}^{T}$, where $W\in R^{1*d}$ is the weight matrix of the linear projection layer. Subsequently, we use a 12-layer standard transformer encoder to extract its features. The entire process can be represented as $f\left(m_{HSI}\right)={Enc}_{12}\left(W*m_{HSI}^{T}\right)$.
@@
Table: The table modality consists of several rows and columns, where each row represents a record and each column represents a different attribute, i.e., $m_{Table}\in R^{{row}*{col}}$. The modal encoder for the table modality inherits the design from TabFormer[36]: first, based on the different degrees of discreteness for each column attribute, we use independent embedding layers to encode discrete and continuous values separately. Subsequently, we employ a single-layer transformer encoder to further extract features. The entire process can be formalized as $f\left(m_{Table}\right)={Enc}_{1}\left(Emb(m_{Table})\right)$.
@@
Trajectory: The trajectory modality reflects the changing information of an object over time and space and is composed of a series of two-dimensional coordinate points, i.e., $m_{Trajectory}\in R^{l*2}$, where $l$ represents the sequence length of trajectory points. The encoder for the trajectory modality inherits the design from TUTR[37]: first, a linear layer is used to expand the dimensions of the two-dimensional trajectory features. This step can be formalized as ${W*m}_{Trajectory}$, where $W\in R^{d*\left(l*2\right)}$ is the weight matrix of the linear projection layer. Then, we use a 2-layer transformer encoder to extract its features. The entire process can be formalized as $f\left(m_{Trajectory}\right)={Enc}_{2}\left({W*m}_{Trajectory}\right)$.
@@
SAR: The synthetic aperture radar (SAR) modality is a type of active remote sensing that reflects the electromagnetic characteristics of objects with respect to microwave backscatter. Due to differences in polarization modes, the final product of SAR imagery is typically a two-band or single-band image, i.e., $m_{SAR}\in R^{H*W*2}$. Therefore, we designed a simple three-layer convolutional network as the modal encoder.
@@
Infrared: The dataset we selected for infrared detection was SYSU-MM01, which is, in fact, a multimodal reidentification dataset comprising both visible and near-infrared (NIR) modalities. The model needs to simultaneously process infrared and visible light images, i.e., $m_{Infrared}=\{{I_{r}\in R}^{H*W*3},{I_{i}\in R}^{H*W*3}\}$. We inherited and slightly modified the design of DDAG[38], discarding the parameters of the shared ResNet-50 layers in DDAG and employing two independently parameterized ResNet-50 networks. These networks extract features from infrared and visible light images, and this process can be formalized as $f\left(m_{Infrared}\right)={Enc}_{ResNet}\left(I_{r}\right)\oplus\ {Enc}_{ResNet}\left(I_{i}\right)$.
@@
Graph: A graph is constructed from a series of nodes and edges, where the attributes of the nodes and the adjacency attributes of the nodes reflect the majority of the features of the graph, i.e., $m_{Graph}\in R^{K*d}$, where K is the number of nodes and d is the node feature dimension. In AllSpark, the modality encoder of a graph is based on the STAEformer[39], whose main design idea is to first use a linear layer to extend its feature dimension and then use several embedding layers to separately encode features such as the node’s characteristics, spatial characteristics, and temporal characteristics. The entire process can be formalized as $f\left(m_{Graph}\right)={Emb}_{node}\left({W*m}_{Graph}^{T}\right)\oplus{Emb}_{spatial}\left({W*m}_{Graph}^{T}\right)\oplus{Emb}_{time}\left({W*m}_{Graph}^{T}\right)$, where $W\in R^{{hidden}*d}$ is the weight of the linear layer.
@@
Oblique Photography: The oblique photography modality consists of images from multiple views, i.e., $m_{Oblique}=\{V_{1}\in R^{C*H*W},\ V_{3}\in R^{C*H*W},V_{3}\in R^{C*H*W},\ldots\ ,V_{n}\in R^{C*H*W}\}$, where $V_{i}$ represents the image of each view and n is the number of views. In the encoder for the oblique photography modality, we use the standard ViT model to extract features from multiple view images and concatenate them. The process can be formalized as $f\left(m_{Oblique}\right)={Enc}_{12}\left(V_{1}\right)\oplus{Enc}_{12}\left(V_{2}\right)\oplus{Enc}_{12}\left(V_{3}\right)\oplus\ldots\oplus{Enc}_{12}\left(V_{n}\right)$.
III-C3 3D modal
@@
Video: The video modality can be considered as the stacking of images in the temporal dimension, i.e., $m_{video}\in R^{T*C*H*W}$. The significant characteristic of the video modality is its high redundancy in both temporal and spatial dimensions. To address this issue, the encoder for the video modality in AllSpark is based on TubeViT[40], which performs sparse sampling on highly redundant data in both time and space to reduce redundancy. The features obtained from sparse sampling are subsequently fed into a 6-layer transformer encoder for further feature extraction. The process can be formalized as $f\left(m_{video}\right)={Enc}_{6}\left({Sparse}(m_{video})\right)$.
@@
PointClouds: PointClouds are typically composed of three-dimensional coordinates and feature values $m_{PointCloud}\in R^{K*\left(d+3\right)}$, where K represents the number of three-dimensional points and d represents the dimensionality of the point cloud features, reflecting information such as the spatial position, shape, colour, and texture of objects. The encoder for the point cloud modality inherits the design from PointBERT[41]: first, point cloud data are grouped and encoded to unify the number of points simultaneously inputted. This step can be represented as $PointGroup=Grouper\left(m_{PointCloud}\right)\in R^{G*N*3}$, where $G$ represents the number of groups and $N$ represents the number of points in each group. Next, the grouped results are input into a one-dimensional convolutional layer to extract feature vectors for each group:$f_{Group}=Conv1d\left(PointGroup\right)\in R^{G*d}$. Finally, the feature vectors for each group are input into a standard 12-layer transformer encoder to extract their global features. The entire process can be formalized as $f\left(m_{PointCloud}\right)=Enc_{12}\left(Conv1d(Grouper(m_{PointCloud}))\right)$.
III-D Modal2Language Bridge